OpenStudy (anonymous):

The sets of numbers 3, 4, 5 and 8, 15, 17 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (loser66):

I forgot what Pythagorean Theorem says. Do you remember?

OpenStudy (anonymous):

it says a^2 + b^2 =c^2

OpenStudy (loser66):

oh, so, for 3 , 4,5 we can check, right? 3^2 = 9 4^2 = 16 5^2 =25 and we have 25 = 16 +9 ahha!!! bingo we have c^2 = a^2 + b^2 where a = 4, b=3 and c =5, right???

OpenStudy (loser66):

Am I right? are you with me?? @otaku789

OpenStudy (anonymous):

They are triples because they follow the rule of \( a^2 + b^2 = c^2 \) and it forms a right triangle and they are positive integers. \( a^2 + b^2 = c^2 \) \( 3^2 + 4^2 = 5^2 \) \( 9 + 16 = 25 \) \( 25= 25 \) is true so it makes 3,4 and 5 triples. Same for the other set of integers.